A G.p has 8terms.its first and last terms are 0.3 and 38.4.calculate
commom ratio

Answer by Boreal(15235)

term a1=0.3
common ratio r
term a2=0.3*r
term a8=0.3*r^7=38.4
therefore, r^7=38.4/0.3=128
r^7=128
r=2
the terms are 0.3, 0.6, 1.2, 2.4, 4.8, 9.6, 19.2, and 38.4